OpenStudy (anonymous):

Write a two-step equation that needs the steps of addition then division in order to be solved, and then solve it. Show your work.

OpenStudy (anonymous):

Do you know the opposite of the operations of addition and division?

OpenStudy (anonymous):

Addition (+) is an operation that combines two numbers, variables, or expressions together (e.g., 5+3) Division (/) is an operation that takes a number, variable, or expression and splits it into groups of the divisor (number in the denominator) (e.g., 6/3)<---- in that case the 3 is the divisor

OpenStudy (anonymous):

2x - 3 = 7 -- add 3 to both sides 2x - 3 + 3 = 7 + 3 2x = 10 -- divide both sides by 2 2x/2 = 10/2 x = 5

Miracrown (miracrown):